When set at the standard position, Autopitch can throw hard balls toward a batter at an average speed of 60 mph. Autopitch devices are made for both major- and minor-league teams to help them improve their batting averages. Autopitch executives take samples of 10 Autopitch devices at a time to monitor these devices and to maintain the highest quality. The average range is 3 mph. Using control-chart techniques; determine control-chart limits for averages and ranges for Autopitch.
Answer #1

From the available information, n 10, 60, R 3 At n 10. factors for computing control chart limits table gives, A 0.308, D D,-

Control chart limits for averages: 60 (0.308)(3) = 60 + 0.924 60.924 LCL =x-AR -60-(0.308)(3) -60 0.924 = 59.076 Therefore, t